Events A and B are mutually exclusive. Suppose event A occurs with probability 0.59 and event...

Events

A

and

B

are mutually exclusive. Suppose event

A

occurs with probability

0.59

and event

B

occurs with probability

0.38

.

  1. Compute the probability that

    A

    occurs or

    B

    does not occur (or both).
  2. Compute the probability that either

    A

    occurs without

    B

    occurring or

    B

    occurs without

    A

    occurring.
Answer #1

We are given here that A and B are mutually exclusive here. Therefore P(A and B) = 0

Also, we are given here that: A occurs with probability 0.59 and event B occurs with probability 0.38.
Therefore, P(A) = 0.59 and P(B) = 0.38

a) The probability here is computed as:
P(A occurs or B does not occur) = 1 - P(B occurs and A does not occur) = 1 - P(B) = 1 - 0.38 = 0.62

Therefore 0.62 is the required probability here.

b) Probability that either A occurs without B occurring or B occurs without A occurring is computed here as:
= P(A) + P(B)
= 0.59 + 0.38 = 0.97

Therefore 0.97 is the required probability here.